# Langya: Revolutionizing Cross-Spatiotemporal Ocean Forecastin

### v1

@@ 1 Main
Ocean forecasting aims to forecast future ocean dynamic states and is crucial for understanding ocean activities. Numerical global ocean forecasting systems (GOFSs), such as the Mercator Ocean physical system (PSY4) and the Real-Time Ocean Forecasting System (RTOFS), rely on geophysical fluid dynamics and partial differential equations (PDEs) to describe future ocean conditions, and such systems remain widely used worldwide. However, improving the inference accuracy and speed of numerical forecast models is a significant challenge. First, numerical forecasting heavily depends on parameterizations and the extent of human knowledge regarding oceanic physical laws. Approximation functions can exacerbate forecast errors. In addition, even on supercomputers equipped with hundreds of computing nodes, it takes several hours to generate a single forecast. Large AI models are reshaping interactions between humans and technology through their massive learning parameters and powerful learning capabilities, e.g., ChatGPT and Sora. Such advances open the way for the development of geophysical system forecasting, including large AI models for weather (e.g., FourcastNet, PanGu-Weather) and ocean (e.g., XiHe [1], AI-GOMS [2]) forecasts. These large models are mainly data-driven, recasting air state variables (ASVs) and ocean state variables (OSVs) forecasting problems as autoregressive video generation, thereby avoiding much computational complexity and uncertainties inherent in physical processes. Moreover, specialized computing devices such as graphics processing units (GPUs) offer extremely fast inference speeds. By investigating and practicing ocean forecasting, we summarize two key barriers in current large ocean forecasting models: (1) Future OSVs are forecasted autoregressively in a step-by-step manner, with temporal information not incorporated into the training and inference processes. It neglects the time-dependent evolution of OSV fields and their interdependencies, making it impossible to achieve cross-timescale forecasts. (2) Air and ocean data are concatenated as model inputs to represent air-sea interactions, a simple way that cannot effectively simulate these interactions nor integrate their underlying mechanisms into the training and inference processes. Although researchers agree that large AI models hold tremendous potential for ocean forecasting, it should be acknowledged that revolutionary breakthroughs are still needed before an AI ocean forecasting system capable of cross-timescale, air-sea coupled forecasts can be realized. Fortunately, the breakthroughs have now come as expected. Here, we present the LangYa forecasting system (see the Methods section for "LangYa"), a powerful large AI-based model for cross-spatiotemporal, air-sea coupled ocean forecasting. LangYa delivers faster and more accurate forecasts for OSVs (temperature, salinity, and currents) than both traditional numerical weather prediction (NWP) systems and XiHe (an influential large AI-based model). Our contributions are fourfold: (1) We propose a large-language-model-based (LLM-based) Time Embedding Module that achieves high spatiotemporal resolution (1/12°, 1 day) forecasts for OSVs. It enables cross-timescale forecasts (up to 7 days) without iterative steps, allowing training and inference for future 7-day OSVs with one-off. (2) We develop an asynchronous cross-iterative random sampling strategy to simulate the atmospheric’s stochastic interaction with the ocean, improving the forecasting skill for variables with different rates of change, such as sea temperature and currents. (3) We introduce an ocean self-attention module based on cosine attention, enabling more stable model training and faster convergence. With this module, a single epoch of distributed training on global OSVs from 1993 to 2019 can be completed in just 0.5 days. (4) We propose an adaptive loss function for thermocline, which significantly improves the accuracy of forecasts at a 100 m depth thermocline, yielding about a 50% improvement in flow forecast accuracy compared to other methods. Using GLORYS12 reanalysis data, we verify that LangYa can achieve stable training and convergence, and effectively yield real-time, accurate, multivariate forecasts across spatiotemporal scales, especially for the thermocline layer.
@@
2 Cross-SpatioTemporal and Air–Sea Coupled Ocean Forecasting System
@@
LangYa adopts the idea of data-driven that builds ocean forecasting as a video prediction task. The design contains training and testing phases by using oceanic and atmospheric reanalysis data as inputs and generating forecasts for Ocean State Variables (OSVs) — such as sea temperature, salinity, and currents — through single-pass inference over 1–7 days. The system uses 11,322 samples and mitigates overfitting risks by randomly shuffling sample order at the beginning of each training iteration. The training was conducted using a distributed data-parallel (DDP) strategy on a cluster of 16 NVIDIA A800 GPUs, completed in just 14 days. The core architecture of the LangYa system is shown in Figure 1a, which integrates multiple innovative components. The input state variables include ASVs, which are u- and v-components of wind speed, temperature, and relative humidity across 1000 to 200 hPa. Sea Surface State Variables (SSSVs): u- and v-components of 10-m wind speed, sea surface temperature, and mean sea level pressure. OSVs: Temperature, salinity, and currents spanning depths of 0 to -500 m. OSVs, ASVs, and SSSVs derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) fifth-generation reanalysis dataset (ERA5) and GLORYS12 reanalysis datasets. The data is complemented by temporal information processed via a time-embedding module based on LLM (Figure 1b). The module extracts spatial features for individual time steps and temporal features across multiple time steps, and then integrates prior knowledge from the LLM to encode forecast durations. The unique fusion of spatiotemporal data supports highly efficient cross-spatiotemporal forecasting. A key innovation of LangYa is its Asynchronous Cross-Iterative Random Sampling Strategy (Figure 1c), which adaptively links ASVs, SSSVs, and OSVs to simulate air-sea interactions. Unlike traditional methods that rely on physics-based assumptions or simple data concatenation, this strategy learns spatiotemporal relationships and effectively captures air–sea variation patterns, significantly enhancing forecast accuracy without introducing additional physical constraints. The processed input data are passed through an encoder-decoder architecture equipped with an Ocean Self-Attention Module (Figure 1d). The module, based on cosine attention, ensures stable training and accelerates convergence by reducing gradient explosions common in high-dimensional data processing. Additionally, LangYa incorporates an Adaptive Thermocline Loss Function (Figure 1e) to address the challenges of forecasting thermocline mutations. The loss function is tailored to capture the steep vertical gradients and dynamic variability of the thermocline, improving forecast accuracy for this critical oceanic layer. By integrating these innovative components, LangYa has achieved several breakthroughs: It significantly reduces temporal distribution discrepancies and simplifies the complexity of air-sea coupling. It supports efficient, accurate, and stable forecasting across spatiotemporal scales. It yields state-of-the-art performance, particularly in challenging regions like thermoclines and areas with intense air-sea interactions. LangYa’s comprehensive design represents a major leap forward in AI-driven ocean forecasting, offering unprecedented efficiency and reliability for global ocean research and operational applications.
@@ 3 Experimental Setting and Main Results
The study trained the LangYa forecasting system using reanalysis data (GLORYS12 and ERA5) and evaluated it with both reanalysis (GLORYS12) and observational datasets. The observational datasets are derived from the GODAE Ocean View Inter-comparison and Validation Task Team (IV-TT) Class 4 framework, which provides observation datasets from Argo, Jason-1, Jason-2, and Envisat CLS AVISO level 3 satellite altimeters. 27 years of GLORYS12 data (1993–2019) were used for training LangYa, while data from 2020–2021 were reserved for testing. LangYa forecasts 128 variables, including temperature, salinity, and zonal and meridional ocean currents across 32 depth layers. Compared with reanalysis and observational data, LangYa achieved lower RMSEs for temperature and ocean currents (both zonal and meridional) than XiHe, XiHe-Autoregression (XiHe-AR), PSY4, BLK, GIOPS, and FOAM numerical models (Figure 2). Notably, the released XiHe version consists of 20 separate models, each designed to forecast OSVs for 1–10 days in either the upper or lower ocean layers. XiHe-AR refers to an autoregressive approach that generates multiday forecasts by sequentially chaining two 1-day XiHe models, resulting in higher cumulative errors. For salinity, LangYa’s forecast accuracy was comparable to XiHe and the numerical models but outperformed XiHe-AR.
Figure 2: Performance Comparison of LangYa, XiHe, XiHe-AR, and Numerical Models Across OSVs. A: RMSE of the u-component of ocean currents for 1–7 day forecasts. B: RMSE of the v-component of ocean currents for 1-7 day forecasts. C: RMSE of the ocean temperature for 1–7 day forecasts. D: RMSE of the salinity for 1–7 Day Forecasts.
In addition, LangYa achieved an inference time of 1 second for forecasts of any scale (1–7 days) on a single GPU, which is over 10,000 times faster than operational numerical models. Results from the 2020–2021 test data indicate that LangYa provided "perfect" quantitative forecasts (RMSE), particularly for challenging ocean phenomena such as thermoclines. Compared to the world’s leading OSVs forecasting systems (e.g., XiHe, PSY4, BLK, GIOPS, and FOAM numerical models), LangYa significantly reduced cumulative forecast errors and demonstrated remarkable improvements in thermocline forecasting.
@@
Considering the significant differences in seawater characteristics across ocean basins, the global ocean was divided into eight basins (Northwest Pacific, Northeast Pacific, North Atlantic, North Indian Ocean, South Atlantic, South Pacific, South Indian Ocean, and Arctic Ocean) to evaluate LangYa’s forecasting performance. Figure 4 presents LangYa’s forecasting accuracy for four OSVs across these eight basins over 1–7 days. The results show substantial variations in RMSEs among the different ocean basins for forecasting ocean variables. Specifically, the Arctic Ocean exhibited the lowest RMSEs for U, V, and T, indicating that LangYa’s performance in this region is the most stable and accurate. This is due to the relatively stable hydrodynamic conditions and fewer drastic changes in the Arctic. Conversely, forecasting OSVs in the North Indian Ocean presented significant challenges, as monsoons, freshwater input, and complex oceanic dynamics influence the North Indian basin. The intricate physical processes and higher uncertainties posed difficulties for LangYa’s forecasts. In contrast, LangYa demonstrated more consistent forecasting performance in the other basins, showcasing the robustness of the LangYa forecasting system.
@@ 4 Thermocline Forecast
The thermocline, a transitional layer in the ocean characterized by a sharp temperature gradient, is a critical interface for energy exchange, water mass mixing, and biogeochemical processes. Typically located at depths of 100-200 m, the thermocline exhibits complex regional and seasonal variations due to its involvement in multiple coupled physical processes, such as ocean stratification, vertical mixing, and air-sea interactions. Accurate forecasting of the thermocline is essential for advancing the understanding of these processes and their implications for global climate systems, yet it has long been a challenge due to the steep gradients and non-linear dynamics in this layer.
@@ 5 Discussion
Despite its achievements, LangYa faces several limitations in practical applications: (1) LangYa is trained and tested on reanalysis data, whereas operational forecasting systems primarily rely on numerical assimilation data. The differences between reanalysis and numerical assimilation data may reduce model performance in applications, requiring further research into cross-data-source adaptability. (2) LangYa currently supports short-term forecasts of 1–7 days and has not yet been extended to medium- and long-term forecasts (e.g., monthly or seasonal scales). Addressing this challenge would require advancements in multi-scale coupling and capturing long-term dynamic changes. (3) LangYa focuses on forecasting OSVs and has not yet been extended to directly forecast ocean phenomena such as tropical cyclone tracks or El Niño events. The ocean phenomena hold greater scientific and practical value but require further breakthroughs in multivariable collaborative modeling. In the future, the integration of LangYa with numerical models will represent a novel approach to ocean forecasting research. The synergy will enable intelligent forecasting to expand from single-variable predictions to complex phenomena, further enhancing the model’s applicability and scientific rigor.

### v2

@@ 1 Main
Ocean forecasting aims to forecast future ocean dynamic states and is crucial for understanding ocean activities. Numerical global ocean forecasting systems (GOFSs), such as the Mercator Ocean physical system (PSY4) [1] and the Real-Time Ocean Forecasting System (RTOFS) [2], rely on geophysical fluid dynamics and partial differential equations (PDEs) to describe future ocean conditions, and such systems remain widely used worldwide. However, improving the inference accuracy and speed of numerical forecast models is a significant challenge. First, numerical forecasting heavily depends on parameterizations and the extent of human knowledge regarding oceanic physical laws. Approximation functions can exacerbate forecast errors. In addition, even on supercomputers equipped with hundreds of computing nodes, it takes several hours to generate a single forecast. Large AI models are reshaping interactions between humans and technology through their massive learning parameters and powerful learning capabilities, e.g., ChatGPT and Sora. Such advances open the way for the development of geophysical system forecasting, including large AI models for weather (e.g., FourcastNet, PanGu-Weather) and ocean (e.g., XiHe [3], AI-GOMS [4]) forecasts. These large models are mainly data-driven, recasting air state variables (ASVs) and ocean state variables (OSVs) forecasting problems as autoregressive video generation, thereby avoiding much computational complexity and uncertainties inherent in physical processes. Moreover, specialized computing devices such as graphics processing units (GPUs) offer extremely fast inference speeds. By investigating and practicing ocean forecasting, we summarize two key barriers in current large ocean forecasting models: (1) Future OSVs are forecasted autoregressively in a step-by-step manner, with temporal information not incorporated into the training and inference processes. It neglects the time-dependent evolution of OSV fields and their interdependencies, making it impossible to achieve cross-timescale forecasts. (2) Air and ocean data are concatenated as model inputs to represent air-sea interactions, a simple way that cannot effectively simulate these interactions nor integrate their underlying mechanisms into the training and inference processes. Although researchers agree that large AI models hold tremendous potential for ocean forecasting, it should be acknowledged that revolutionary breakthroughs are still needed before an AI ocean forecasting system capable of cross-timescale, air-sea coupled forecasts can be realized. Fortunately, the breakthroughs have now come as expected. Here, we present the LangYa forecasting system (see the Methods section for "LangYa"), a powerful large AI-based model for cross-spatiotemporal, air-sea coupled ocean forecasting. LangYa delivers faster and more accurate forecasts for OSVs (temperature, salinity, and currents) than both traditional numerical weather prediction (NWP) systems and XiHe (an influential large AI-based model). Our contributions are fourfold: (1) We propose a large-language-model-based (LLM-based) Time Embedding Module that achieves high spatiotemporal resolution (1/12°, 1 day) forecasts for OSVs. It enables cross-timescale forecasts (up to 7 days) without iterative steps, allowing training and inference for future 7-day OSVs with one-off. (2) We develop an asynchronous cross-iterative random sampling strategy to simulate the atmospheric’s stochastic interaction with the ocean, improving the forecasting skill for variables with different rates of change, such as sea temperature and currents. (3) We introduce an ocean self-attention module based on cosine attention, enabling more stable model training and faster convergence. With this module, a single epoch of distributed training on global OSVs from 1993 to 2019 can be completed in just 0.5 days. (4) We propose an adaptive loss function for thermocline, which significantly improves the accuracy of forecasts at a 100 m depth thermocline, yielding about a 50% improvement in flow forecast accuracy compared to other methods. Using GLORYS12 reanalysis data, we verify that LangYa can achieve stable training and convergence, and effectively yield real-time, accurate, multivariate forecasts across spatiotemporal scales, especially for the thermocline layer.
@@
2 Cross-SpatioTemporal and Air-Sea Coupled Ocean Forecasting System
@@
LangYa adopts the idea of data-driven that builds ocean forecasting as a video prediction task. The design contains training and testing phases by using oceanic and atmospheric reanalysis data as inputs and generating forecasts for Ocean State Variables (OSVs) — such as sea temperature, salinity, and currents — through single-pass inference over 1-7 days. The system uses 11,322 samples and mitigates overfitting risks by randomly shuffling sample order at the beginning of each training iteration. The training was conducted using a distributed data-parallel (DDP) strategy on a cluster of 16 NVIDIA A800 GPUs, completed in just 14 days. The core architecture of the LangYa system is shown in Figure 1a, which integrates multiple innovative components. The input state variables include ASVs, which are u- and v-components of wind speed, temperature, and relative humidity across 1000 to 200 hPa. Sea Surface State Variables (SSSVs): u- and v-components of 10-m wind speed, sea surface temperature, and mean sea level pressure. OSVs: Temperature, salinity, and currents spanning depths of 0 to -500 m. OSVs, ASVs, and SSSVs derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) fifth-generation reanalysis dataset (ERA5) and GLORYS12 reanalysis datasets. The data is complemented by temporal information processed via a time-embedding module based on LLM (Figure 1b). The module extracts spatial features for individual time steps and temporal features across multiple time steps, and then integrates prior knowledge from the LLM to encode forecast durations. The unique fusion of spatiotemporal data supports highly efficient cross-spatiotemporal forecasting. A key innovation of LangYa is its Asynchronous Cross-Iterative Random Sampling Strategy (Figure 1c), which adaptively links ASVs, SSSVs, and OSVs to simulate air-sea interactions. Unlike traditional methods that rely on physics-based assumptions or simple data concatenation, this strategy learns spatiotemporal relationships and effectively captures air-sea variation patterns, significantly enhancing forecast accuracy without introducing additional physical constraints. The processed input data are passed through an encoder-decoder architecture equipped with an Ocean Self-Attention Module (Figure 1d). The module, based on cosine attention, ensures stable training and accelerates convergence by reducing gradient explosions common in high-dimensional data processing. Additionally, LangYa incorporates an Adaptive Thermocline Loss Function (Figure 1e) to address the challenges of forecasting thermocline mutations. The loss function is tailored to capture the steep vertical gradients and dynamic variability of the thermocline, improving forecast accuracy for this critical oceanic layer. By integrating these innovative components, LangYa has achieved several breakthroughs: It significantly reduces temporal distribution discrepancies and simplifies the complexity of air-sea coupling. It supports efficient, accurate, and stable forecasting across spatiotemporal scales. It yields state-of-the-art performance, particularly in challenging regions like thermoclines and areas with intense air-sea interactions. LangYa’s comprehensive design represents a major leap forward in AI-driven ocean forecasting, offering unprecedented efficiency and reliability for global ocean research and operational applications.
@@ 3 Experimental Setting and Main Results
The study trained the LangYa forecasting system using reanalysis data (GLORYS12 and ERA5) and evaluated it with both reanalysis (GLORYS12) and observational datasets. The observational datasets are derived from the GODAE Ocean View Inter-comparison and Validation Task Team (IV-TT) Class 4 framework, which provides observation datasets from Argo, Jason-1, Jason-2, and Envisat CLS AVISO level 3 satellite altimeters. 27 years of GLORYS12 data (1993-2019) were used for training LangYa, while data from 2020-2021 were reserved for testing. LangYa forecasts 128 variables, including temperature, salinity, and zonal and meridional ocean currents across 32 depth layers. Compared with reanalysis and observational data, LangYa achieved lower RMSEs for temperature and ocean currents (both zonal and meridional) than XiHe, XiHe-Autoregression (XiHe-AR), PSY4 [1], BLUElinK OceanMAPS (BLK) [5], Global Ice Ocean Prediction System (GIOPS) [6], and Forecast Ocean Assimilation Model (FOAM) numerical models (Figure 2). Notably, the released XiHe version consists of 20 separate models, each designed to forecast OSVs for 1-10 days in either the upper or lower ocean layers. XiHe-AR refers to an autoregressive approach that generates multiday forecasts by sequentially chaining two 1-day XiHe models, resulting in higher cumulative errors. For salinity, LangYa’s forecast accuracy was comparable to XiHe and the numerical models but outperformed XiHe-AR.
Figure 2: Performance Comparison of LangYa, XiHe, XiHe-AR, and Numerical Models Across OSVs. A: RMSE of the u-component of ocean currents for 1-7 day forecasts. B: RMSE of the v-component of ocean currents for 1-7 day forecasts. C: RMSE of the ocean temperature for 1-7 day forecasts. D: RMSE of the salinity for 1-7 Day Forecasts.
In addition, LangYa achieved an inference time of 1 second for forecasts of any scale (1-7 days) on a single GPU, which is over 10,000 times faster than operational numerical models. Results from the 2020-2021 test data indicate that LangYa provided "perfect" quantitative forecasts (RMSE), particularly for challenging ocean phenomena such as thermoclines. Compared to the world’s leading OSVs forecasting systems (e.g., XiHe, PSY4, BLK, GIOPS, and FOAM numerical models), LangYa significantly reduced cumulative forecast errors and demonstrated remarkable improvements in thermocline forecasting.
@@
Considering the significant differences in seawater characteristics across ocean basins, the global ocean was divided into eight basins (Northwest Pacific, Northeast Pacific, North Atlantic, North Indian Ocean, South Atlantic, South Pacific, South Indian Ocean, and Arctic Ocean) to evaluate LangYa’s forecasting performance. Figure 4 presents LangYa’s forecasting accuracy for four OSVs across these eight basins over 1-7 days. The results show substantial variations in RMSEs among the different ocean basins for forecasting ocean variables. Specifically, the Arctic Ocean exhibited the lowest RMSEs for U, V, and T, indicating that LangYa’s performance in this region is the most stable and accurate. This is due to the relatively stable hydrodynamic conditions and fewer drastic changes in the Arctic. Conversely, forecasting OSVs in the North Indian Ocean presented significant challenges, as monsoons, freshwater input, and complex oceanic dynamics influence the North Indian basin. The intricate physical processes and higher uncertainties posed difficulties for LangYa’s forecasts. In contrast, LangYa demonstrated more consistent forecasting performance in the other basins, showcasing the robustness of the LangYa forecasting system.
@@ 4 Thermocline Forecast
The thermocline, a transitional layer in the ocean characterized by a sharp temperature gradient, is a critical interface for energy exchange, water mass mixing, and biogeochemical processes. Typically located at depths of 100-200 meter, the thermocline exhibits complex regional and seasonal variations due to its involvement in multiple coupled physical processes, such as ocean stratification, vertical mixing, and air-sea interactions. Accurate forecasting of the thermocline is essential for advancing the understanding of these processes and their implications for global climate systems, yet it has long been a challenge due to the steep gradients and non-linear dynamics in this layer.
@@ 5 Discussion
Despite its achievements, LangYa faces several limitations in practical applications: (1) LangYa is trained and tested on reanalysis data, whereas operational forecasting systems primarily rely on numerical assimilation data. The differences between reanalysis and numerical assimilation data may reduce model performance in applications, requiring further research into cross-data-source adaptability. (2) LangYa currently supports short-term forecasts of 1-7 days and has not yet been extended to medium- and long-term forecasts (e.g., monthly or seasonal scales). Addressing this challenge would require advancements in multi-scale coupling and capturing long-term dynamic changes. (3) LangYa focuses on forecasting OSVs and has not yet been extended to directly forecast ocean phenomena such as tropical cyclone tracks or El Niño events. The ocean phenomena hold greater scientific and practical value but require further breakthroughs in multivariable collaborative modeling. In the future, the integration of LangYa with numerical models will represent a novel approach to ocean forecasting research. The synergy will enable intelligent forecasting to expand from single-variable predictions to complex phenomena, further enhancing the model’s applicability and scientific rigor.